OpenStudy (anonymous):

In humans, the ABO blood type is under the control of autosomal multiple alleles. Color blindness is a recessive X-linked trait. Mom is type A and is color-blind. Dad is type B and has normal vision. They have a daughter who is type O and has normal vision. What is the probability that their next child will be a son who is type AB and is color-blind? make the answer into a fraction

OpenStudy (aaronq):

Since the daughter is type O we know that each parent has to have an O allele. Mom also has both X chromosomes carrying the color-blind allele. Dad has 1 normal X (and a Y). So the parental genotypes are: Mom: AO and \(\sf X^c X^c\) Dad: BO and XY The crosses are then What is the probability that their next child will be a son who is type AB and is color-blind? Probability of having a son =\(\dfrac{1}{2}\) Probability of having a kid with AB = \(\dfrac{1}{4}\) (see first punnet square) Probability of having a son who is color-blind =1 (see second punnett square) (all boys will be colo-blind because the mom can only give the \(X^c\) allele. P(son that is AB and color-blind)=\(\dfrac{1}{2}*\dfrac{1}{4}*1=\dfrac{1}{8}\)